OpenStudy (anonymous):

A cylinder has a radius of 4x + 2 and a height of 5x + 4. Which polynomial in standard form best describes the total volume of the cylinder? Use the formula V = r²h for the colume of a cylinder. 80x³ + 144x² + 84x + 16 80x³ – 144x² + 84x + 16 80x³ + 144x² + 84x – 16 80x³ + 144x² – 84x + 16

OpenStudy (anonymous):

the pie shin is after 80 144 84 and 16 on all answers

OpenStudy (asnaseer):

first - do you know what the volume of a cylinder with base radius r and height h is?

OpenStudy (asnaseer):

looks like you've been given the formula for it

OpenStudy (asnaseer):

\[V=\pi r^2h\]so just substitute r=4x+2 and h=5x+4 into this equation.

OpenStudy (asnaseer):

you can do this in steps, first work out the value of \(r^2\), i.e. \((4x+2)^2\)

OpenStudy (anonymous):

80x³ + 144x² + 84x + 16 A cylinder has a radius of 4x + 2 and a height of 5x + 4. V = (pi r²)h = p * (4x + 2 )² (5x + 4) = pi ( 16x^2 + 16x + 4) (5x + 4) = pi ( 80x³ + 80x² + 20x) + ( 64x² + 64x + 16) = pi ( 80x³ + 144x² + 84x + 16)